the sum of twice a number and 7 is 140


Answer by DrBeeee(684)

Let x = the number
Then we have
(1) 2*x + 7 = 140 or
(2) 2*x = 133 or
(3) x = 133/2 or
(4) x = 66.5
Is (2*66.5 + 7 = 140)?
Is (133 + 7 = 140)?
Is (140 = 140)? Yes
Answer: the number is 66.5